OpenStudy (anonymous):

Medals 0 How will I solve this "A 30% of barium chloride is mixed with 10 liters of water to form 20% solution. How many Liters of original solution did we start with?"

OpenStudy (anonymous):

\[.3x+10=.2(x+10)\] is a start

OpenStudy (anonymous):

probably easier to multiply both sides by 10 and start with \[3x+100=2(x+10)\]

OpenStudy (anonymous):

I dot get it kindly explain:)